About 10% of the population has a particular genetic disorder. 700 people are randomly selected.

Find the mean for the number of people with the genetic disorder in such groups of 700.
Answer #1

Solution :

Given that,

Using binomial distribution,

Mean = \mu = n * p = 700 * 0.70 = 70